# The bracket of the exceptional Lie algebra E8

### v1

@@ 1. Introduction
The exceptional Lie groups and Lie algebras $\mathrm{G}_{2}$, $\mathrm{F}_{4}$, $\mathrm{E}_{6}$, $\mathrm{E}_{7}$, $\mathrm{E}_{8}$ were first described by Wilhelm Killing [10, 4]. The exceptional Lie groups have since made many appearances in geometry, topology and physics. However, to this day they still remain somewhat elusive and mysterious.
@@
Of the five exceptional Lie algebras $\operatorname{\mathfrak{g}}_{2}$, $\operatorname{\mathfrak{f}}_{4}$, $\operatorname{\mathfrak{e}}_{6}$, $\operatorname{\mathfrak{e}}_{7}$, $\operatorname{\mathfrak{e}}_{8}$ the $248$-dimensional algebra $\operatorname{\mathfrak{e}}_{8}$ is the one with highest dimension, and in many regards also the most difficult to deal with. For instance, since the lowest dimensional non-trivial representation is $248$-dimensional, it is not an option to use a matrix representation of $\operatorname{\mathfrak{e}}_{8}$, an approach that is viable for $\operatorname{\mathfrak{g}}_{2}$, which has a $7$-dimensional faithful representation. For this and other reasons, the title of the article [8] refers to $\mathrm{E}_{8}$ as the “most exceptional Lie group” and it has been argued in [8], that $\mathrm{E}_{8}$ could be called “the Monster of Lie theory”. The difficulties that arise when computations with these objects are required, have led to computer-based approaches as in [13], where structure constants of $\operatorname{\mathfrak{e}}_{8}$ are determined.
While several descriptions and constructions of $\operatorname{\mathfrak{e}}_{8}$ can be found in the literature, see [1, 5, 14], so far, no explicit formula for its bracket appears to be known. In this paper, we give an explicit formula for the Lie bracket of the compact real form of $\operatorname{\mathfrak{e}}_{8}$, which drastically simplifies certain calculations with $\operatorname{\mathfrak{e}}_{8}$. Besides standard linear algebra operations, our formula (4.1) involves multiplication of Cayley numbers and the triality automorphism of $\operatorname{\operatorname{\mathfrak{so}}}(8)$. Our description is based on the decomposition
@@
which is called the Barton-Sudbery description of $\operatorname{\mathfrak{e}}_{8}$, cf. [3], by Baez [2]. In his article [2], Baez writes: “To emphasize the importance of triality, we can rewrite the Barton-Sudbery description of $\operatorname{\mathfrak{e}}_{8}$ as
@@
$\operatorname{\mathfrak{f}}_{4}\cong\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus V_{8}\oplus S^{+}_{8}\oplus S^{-}_{8}.\hbox{''}$ ”
In the article [11], the author obtained an explicit bracket formula (2.5) for $\operatorname{\mathfrak{f}}_{4}$ in order to study a geometric problem on the Cayley hyperbolic plane. This formula is based on the above-mentioned triality description of $\operatorname{\mathfrak{f}}_{4}$. It turns out that there is a natural generalization (4.1) of this formula for the bracket of $\operatorname{\mathfrak{e}}_{8}$. We obtain it by replacing $\operatorname{\operatorname{\mathfrak{so}}}(8)$ with $\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)$ and replacing the octonions $\mathbb{O}$ with the real algebra $\mathbb{O}\otimes\mathbb{O}$, the oct-octonions. The purpose of this paper is to prove that the formula (4.1) actually defines a Lie algebra, i.e., satisfies the Jacobi identity and this Lie algebra can be endowed with an invariant scalar product, showing that the resulting simple real Lie algebra is isomorphic to the compact real form of $\operatorname{\mathfrak{e}}_{8}$. This can also be regarded as a new existence proof for $\operatorname{\mathfrak{e}}_{8}$.
@@ 2. Preliminaries
We identify the real numbers $\mathbb{R}$ with the subalgebra of $\mathbb{O}$ spanned by $e$ and define the real part of an octonion $x$ as the real number $\mathrm{Re}(x):=\frac{1}{2}(x+\bar{x})$ and the pure part as the octonion $\operatorname{Pu}(x):=\frac{1}{2}(x-\bar{x})$. The norm of an octonion $a$ is defined as $|a|=\sqrt{a\bar{a}}$. Recall that $\mathbb{O}$ is a normed algebra, i.e. we have $|ab|=|a|\,|b|$ for all $a,b\in\mathbb{O}$. For $u,v\in\mathbb{R}^{8}=\mathbb{O}$, we denote by $\langle u,v\rangle=\mathrm{Re}(u\bar{v})=u^{t}v$ the standard scalar product of $\mathbb{R}^{8}$.
@@
see [7], [12, §2, Thm. 2]. In [11] the following explicit expression was obtained for the Lie bracket of the simple compact Lie algebra $\operatorname{\mathfrak{f}}_{4}$.
@@ 4. The Lie algebra $\mathfrak{e}_{8}$
It follows from the principle of triality, see [12, §2-3], that
@@
$[\xi,[\eta,\zeta]]+[\eta,[\zeta,\xi]]+[\zeta,[\xi,\eta]]=0$ (4.3)
@@
    where we have used the infinitesimal principle of triality, see [12, §2, Thm. 1], which implies $(Au)v+u\lambda(A)v=\kappa\circ\lambda^{2}(A)(uv)$ for all $A\in\operatorname{\operatorname{\mathfrak{so}}}(8)$ and all $u,v\in\mathbb{O}$.
@@
      00 0 $=e\wedge\bar{t}+\lambda^{2}(\bar{t}\wedge e)+\lambda(\bar{t}\wedge e)=e\wedge\bar{u}+\lambda^{2}(\bar{u}\wedge e)+\lambda(\bar{u}\wedge e).$
@@
      It can be directly verified that $2t\wedge e=R_{t}+L_{t}$ holds for $t\in\operatorname{Pu}(\mathbb{O})$, see [11, Equation (2.2)].
@@
      To see that the term in the last line is zero, one may use the identification (3.1) to write the summands as matrices. Then $-p\otimes q+\bar{p}\otimes\bar{q}$ is a real $8\times 8$ matrix which has non-zero entries only in the first row and the first column and so is $2(\langle e,q\rangle\operatorname{Pu}(p)\otimes e+\langle e,p\rangle e\otimes\operatorname{Pu}(q))$; it is easy to see that these two matrices are equal.
@@ 5. The noncompact simple Lie algebra $\mathfrak{e}_{8(8)}$
Let $\mathfrak{g}=\mathfrak{e}_{8}$, the compact form. It is easy to see from the bracket formula (4.1) and Lemma 3.1 that the map $(A,u,v,w)\mapsto(A,u,-v,-w)$ is an involutive automorphism of $\mathfrak{g}$, corresponding to the Riemannian symmetric space ${\rm E}_{8}/\operatorname{Spin}(16)$. Let $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ be the decomposition into the $(+1)$-eigenspace and $(-1)$-eigenspace. It follows that $\mathfrak{g}=\mathfrak{k}\oplus i\mathfrak{p}\subset\mathfrak{g}\otimes\mathbb{C}$ is the Lie algebra of the isometry group of the dual symmetric space ${\rm E}_{8(8)}/\operatorname{Spin}(16)$, see [9]. Its bracket is given by the above formula. ∎

### v2

@@ 1. Introduction
The exceptional Lie groups and Lie algebras $\mathrm{G}_{2}$, $\mathrm{F}_{4}$, $\mathrm{E}_{6}$, $\mathrm{E}_{7}$, $\mathrm{E}_{8}$ were first described by Wilhelm Killing [13, 6]. The exceptional Lie groups have since made many appearances in geometry, topology, and physics. However, to this day, they still remain somewhat elusive and mysterious.
@@
Of the five exceptional Lie algebras $\operatorname{\mathfrak{g}}_{2}$, $\operatorname{\mathfrak{f}}_{4}$, $\operatorname{\mathfrak{e}}_{6}$, $\operatorname{\mathfrak{e}}_{7}$, $\operatorname{\mathfrak{e}}_{8}$ the $248$-dimensional algebra $\operatorname{\mathfrak{e}}_{8}$ is the one with the highest dimension, and in many regards also the most difficult to deal with. For instance, since the lowest dimensional non-trivial representation is $248$-dimensional, it is not a good option to use a matrix representation of $\operatorname{\mathfrak{e}}_{8}$, an approach that is viable for $\operatorname{\mathfrak{g}}_{2}$, which has a $7$-dimensional faithful representation. For this and other reasons, the title of the article [11] refers to $\mathrm{E}_{8}$ as the “most exceptional Lie group” and it has been argued in [11], that $\mathrm{E}_{8}$ could be called “the Monster of Lie theory”.
Many different descriptions and constructions of $\operatorname{\mathfrak{e}}_{8}$ and other exceptional Lie algebras can already be found in the literature, as will be discussed below. It appears that exceptional Lie algebras and exceptional Lie groups have many different facets and thus there is not a single one-size-fits-all approach to them, but different aspects of exceptional Lie groups lead to different constructions. Let us briefly review some known constructions of $\operatorname{\mathfrak{e}}_{8}$ and other Lie algebras, without aiming for completeness.
Construction using a subalgebra.
One may construct a model for $\operatorname{\mathfrak{g}}=\operatorname{\mathfrak{e}}_{8}$ and other Lie algebras by starting with a subalgebra $\operatorname{\mathfrak{h}}$ of $\operatorname{\mathfrak{g}}$. When restricted to $\operatorname{\mathfrak{h}}$, the adjoint representation of $\operatorname{\mathfrak{g}}$ splits as $\operatorname{\mathfrak{g}}=\operatorname{\mathfrak{h}}\oplus V$, where the real vector space $V$ is an adjoint representation of. In order to define the Lie bracket on $\operatorname{\mathfrak{g}}$, one only has to define the bracket operation on $V\times V$, since the bracket on $\operatorname{\mathfrak{h}}\times\operatorname{\mathfrak{h}}$ and on $\operatorname{\mathfrak{h}}\times V$ is already given. This approach is employed in the articles [1, 7, 15]. Various choices of for the Lie algebra $\operatorname{\mathfrak{h}}$ lead to different models for $\operatorname{\mathfrak{g}}$. It seems to be most natural and convenient to choose the subalgebra $\operatorname{\mathfrak{h}}$ to be of full rank. In the articles [1] and  [15], the subalgebra $\operatorname{\mathfrak{h}}=\operatorname{\mathfrak{so}}_{16}$ is used. In [7] this procedure is applied to 14 different subalgebras of $\operatorname{\mathfrak{e}}_{8}$, but no explicit expressions for the Lie bracket are given.
Construction using derivations.
It is well known that the compact Lie group $\operatorname{\sf{F}}_{4}$ can be identified with the automorphism group of an Albert algebra and thus its Lie algebra can be constructed as the algebra of derivations of this Albert algebra. The Tits construction [17], see also [8, 18], is an elegant and powerful generalization of this fact which also encompasses all the Lie algebras in Freudenthal’s magic square. In [17] and [18], formulae are given for the brackets of these Lie algebras. It is remarked in [19]: “However, the definition of the Lie bracket given by Vinberg is not straightforward, and involves some quite complicated and unintuitive formulae.”. In this paper, we will give a more natural and simpler explicit formula for the bracket of $\operatorname{\mathfrak{e}}_{8}$.
Construction as commutator algebra of matrix algebra.
A relatively new approach was recently proposed in [19]. It imitates the construction of classical Lie algebras arising from associative algebras using the commutator of matrices as a Lie bracket. However, since the octonions, which are used to define the matrix algebra, are non-associative, one cannot use the usual matrix multiplication and the commutator given by it, but the matrix multiplication has to be modified.
In this paper, we give an explicit formula for the Lie bracket of the compact real form of $\operatorname{\mathfrak{e}}_{8}$, which has the potential to drastically simplify calculations with $\operatorname{\mathfrak{e}}_{8}$ in certain situations. Our approach is based on the first of the three construction methods discussed above, namely, the Lie algebra $\operatorname{\mathfrak{g}}=\operatorname{\mathfrak{e}}_{8}$ is built form its subalgebra $\operatorname{\mathfrak{h}}=\operatorname{\mathfrak{so}}_{8}\oplus\operatorname{\mathfrak{so}}_{8}$. As pointed out above, it only remains to define the bracket of $\operatorname{\mathfrak{g}}=\operatorname{\mathfrak{h}}\oplus V$ on $V\times V$ which we can do using a very a natural map, which is essentially just the multiplication in the algebra $\mathbb{O}\otimes\mathbb{O}$. The appeal of this approach is that $\operatorname{\mathfrak{e}}_{8}$ is built from scratch using little more than octonionic multiplication and the resulting formula (4.1) is very concise, symmetric and intuitive. Indeed, besides standard linear algebra operations, our formula only involves multiplication of Cayley numbers and the triality automorphism of $\operatorname{\operatorname{\mathfrak{so}}}(8)$ (which is also described using octonionic multiplication).
Our description is based on the decomposition
@@
which is called the Barton-Sudbery description of $\operatorname{\mathfrak{e}}_{8}$, see [4], by Baez [2].
In [3], Baez writes: “To emphasize the importance of triality, we can rewrite the Barton-Sudbery description of $\operatorname{\mathfrak{e}}_{8}$ as
@@
$\operatorname{\mathfrak{f}}_{4}\cong\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus V_{8}\oplus S^{+}_{8}\oplus S^{-}_{8}.\hbox{''}$
In the article [14], the author obtained an explicit bracket formula (2.5) for $\operatorname{\mathfrak{f}}_{4}$ in order to study a geometric problem on the Cayley hyperbolic plane. This formula is based on the above-mentioned triality description of $\operatorname{\mathfrak{f}}_{4}$. It appears that formulae of this type are particularly suitable for studying geometric problems involving Lie group actions, see also [5], where the formula (2.5) is successfully applied to a geometric problem.
It turns out that there is a natural generalization (4.1) of this formula for the bracket of $\operatorname{\mathfrak{e}}_{8}$. We obtain it by replacing $\operatorname{\operatorname{\mathfrak{so}}}(8)$ with $\operatorname{\operatorname{\mathfrak{so}}}(8)\oplus\operatorname{\operatorname{\mathfrak{so}}}(8)$ and replacing the octonions $\mathbb{O}$ with the real algebra $\mathbb{O}\otimes\mathbb{O}$, the oct-octonions. The purpose of this paper is to prove that the formula (4.1) actually defines a Lie algebra, i.e., satisfies the Jacobi identity and this Lie algebra can be endowed with an invariant scalar product, showing that the resulting simple real Lie algebra is isomorphic to the compact real form of $\operatorname{\mathfrak{e}}_{8}$. This can also be regarded as an independent new existence proof for $\operatorname{\mathfrak{e}}_{8}$.
@@ 2. Preliminaries
We identify the real numbers $\mathbb{R}$ with the subalgebra of $\mathbb{O}$ spanned by $e$ and define the real part of an octonion $x$ as the real number $\mathrm{Re}(x):=\frac{1}{2}(x+\bar{x})$ and the pure part as the octonion $\operatorname{Pu}(x):=\frac{1}{2}(x-\bar{x})$. The norm of an octonion $a$ is defined as $|a|=\sqrt{a\bar{a}}$. Recall that $\mathbb{O}$ is a normed algebra, i.e. we have |ab|=|a||b||ab|=|a|\,|b| for all $a,b\in\mathbb{O}$. For $u,v\in\mathbb{R}^{8}=\mathbb{O}$, we denote by $\langle u,v\rangle=\mathrm{Re}(u\bar{v})=u^{t}v$ the standard scalar product of $\mathbb{R}^{8}$.
@@
see [10], [16, §2, Thm. 2]. In [14] the following explicit expression was obtained for the Lie bracket of the simple compact Lie algebra $\operatorname{\mathfrak{f}}_{4}$.
@@ 4. The Lie algebra $\mathfrak{e}_{8}$
It follows from the principle of triality, see [16, §2-3], that
@@
$\xi,[\eta,\zeta]]+[\eta,[\zeta,\xi]]+[\zeta,[\xi,\eta]]=0$ (4.3)
@@
    where we have used the infinitesimal principle of triality, see [16, §2, Thm. 1], which implies $(Au)v+u\lambda(A)v=\kappa\circ\lambda^{2}(A)(uv)$ for all $A\in\operatorname{\operatorname{\mathfrak{so}}}(8)$ and all $u,v\in\mathbb{O}$.
@@
      $0$ $=e\wedge\bar{t}+\lambda^{2}(\bar{t}\wedge e)+\lambda(\bar{t}\wedge e)=e\wedge\bar{u}+\lambda^{2}(\bar{u}\wedge e)+\lambda(\bar{u}\wedge e).$
@@
      It can be directly verified that $2t\wedge e=R_{t}+L_{t}$ holds for $t\in\operatorname{Pu}(\mathbb{O})$, see [14, Equation (2.2)].
@@
      To see that the term in the last line is zero, one may use the identification (3.1) to write the summands as matrices. Then $p\otimes q+\bar{p}\otimes\bar{q}$ is a real $8\times 8$ matrix which has non-zero entries only in the first row and the first column and so is $2(\langle e,q\rangle\operatorname{Pu}(p)\otimes e+\langle e,p\rangle e\otimes\operatorname{Pu}(q))$; it is easy to see that these two matrices are equal.
@@ 5. The noncompact simple Lie algebra $\mathfrak{e}_{8(8)}$
Let $\mathfrak{g}=\mathfrak{e}_{8}$, the compact form. It is easy to see from the bracket formula (4.1) and Lemma 3.1 that the map $(A,u,v,w)\mapsto(A,u,-v,-w)$ is an involutive automorphism of $\mathfrak{g}$, corresponding to the Riemannian symmetric space ${\rm E}_{8}/\operatorname{Spin}(16)$. Let $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ be the decomposition into the $(+1)$-eigenspace and (1)(-1)-eigenspace. It follows that $\mathfrak{g}=\mathfrak{k}\oplus i\mathfrak{p}\subset\mathfrak{g}\otimes\mathbb{C}$ is the Lie algebra of the isometry group of the dual symmetric space ${\rm E}_{8(8)}/\operatorname{Spin}(16)$, see [12]. Its bracket is given by the above formula. ∎